OpenStudy (anonymous):

bmp or amistre64...can you check my evaluation of y'(x) =the integral from 0 to x of t * f(t) dt + the integral from x to 1 (t-1)f(t) dt if x = 0

OpenStudy (anonymous):

\[y'(0) = \int\limits_{0}^{0} t * f(t)dt + \int\limits_{0}^{}\]

OpenStudy (anonymous):

\[\int\limits_{0}^{1}(t-1) f(t) dt\]

OpenStudy (anonymous):

= 0 + \[\int\limits_{0}^{1}(t-1) f(t) dt\]

OpenStudy (anonymous):

now what do I do with this last phrase, the integral from 0 to 1 with x as the 0 but there are no x's in the integral?